### v5

# Using iterated local alignment to aggregate trajectory data into a traffic flow map

Abstract
Vehicle trajectories are a promising GNSS (Global Navigation Satellite System) data source to compute multi-scale traffic flow maps ranging from the city/regional level to the road level. The main obstacle is that trajectory data are prone to measurement noise. While this is negligible for city level, large-scale flow aggregation, it poses substantial difficulties for road level, small-scale aggregation. To overcome these difficulties, we introduce innovative local alignment algorithms, where we infer road segments to serve as local reference segments, and proceed to align nearby road segments to them. We deploy these algorithms in an iterative workflow to compute locally aligned flow maps. By applying this workflow to synthetic and empirical trajectories, we verify that our locally aligned flow maps provide high levels of accuracy and spatial resolution of flow aggregation at multiple scales for static and interactive maps.
Keywords: GNSS, GPS, rasterisation, multi-scale, route finding
1 Introduction
One of the fundamental quantities in transport planning is a traffic flow map, i.e., a map of the traffic flow levels on the road segments in a road network (Ortúzar and Willumsen, 2011). Whilst traffic flow maps are a rich source of information about vehicle mobility patterns, they are costly in terms of time and resources to compute for any reasonably sized road network. To alleviate this cost burden, most approaches restrict the spatial resolution of the flow map. One of the most common is to place sensors at fixed locations in the road network, whose results are then visualized as a traffic count map. Another common approach are trip intent/recall questionnaires to inform large-scale properties such as trajectory origins and destinations, whose results are then visualized as a desire line flow map/spider diagram (Tobler, 1987). Both of these are simplified flow maps, since the detailed mobility patterns outside of the sensor locations or origin/destination pairs are not known. These unknown patterns can be inferred from other data sources, such as route assignment models (Evans, 1976). While these model-assigned routes are highly detailed, the trade-off is that they are not guaranteed to correspond closely to empirical mobility patterns.
Trajectory data, derived from GNSS (Global Navigation Satellite Systems), combine the large-scale information of road sensor counts or questionnaires, with the small-scale details of model-assigned trajectories. Due to the prevalence of GNSS-enabled devices, such as vehicle navigation guides and mobile telephones, trajectory data can be acquired with low marginal cost whilst at the same time offering extensive spatial coverage and resolution of empirical mobility patterns (Herrera et al., 2010, Andrienko and Andrienko, 2013). Our motivating example of trajectory data are the 1183 trajectories in EPSG 32632 (WGS 84/UTM zone 32N) projected coordinates, collected from a GNSS-enabled mobile phone, from December 2017 to March 2019 by a single driver in Hannover, Germany, with an overall average sampling rate about 1 point per second (Zourlidou et al., 2022). In Figure 1(a), at the city level, the green circles appear to be aligned to the road network.
Figure 1: Trajectory data in Hannover, Germany. (a) Empirical trajectories (green circles). Klusreide road segment (solid yellow circle). (b) Zoom of trajectories and map matched routes (blue lines), with road network (solid black lines), near Klusreide. (d) Zoom of 2 m $\times$ 2 m rasterised flow aggregation. Colour scale is traffic flow in grid cells.
If we zoom in on these trajectory points around Klusriede (solid yellow circle), a residential street near the A2 autobahn, in Figure 1(b), then we observe that the points deviate from the road network (solid black lines). These noisy trajectories are unsuitable for direct aggregation into a flow map, so the usual first step is to apply a map matching procedure (Quddus et al., 2007, Chao et al., 2020). The map matched routes are the blue lines in Figure 1(b). These follow the road network more closely. Map matching is a highly difficult problem and map matching with exact alignment to the underlying road network in a reasonable time frame is usually not feasible. The approach for fast computations taken by the Valhalla routing engine introduces some approximations which do not lead to exact alignment of empirical trajectories with the road network or to each other (Sghaier, 2019). Nonetheless the misalignment is reduced to less than 5 m, which is, for many purposes, including most notably navigation, indeed a suitable result.
For these inexactly aligned trajectories, applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) leads to an inaccurate flow map. The authors resolve this by rasterising the flow map to a raster grid: and within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell. This rasterised flow map is displayed in Figure 1(c). The colour scale for the flow ranges from 100 (violet) to 300 (pink) to 600 (orange). The 2 m $\times$ 2 m raster grid obscures the road network at this scale. The resolution of a rasterised aggregation flow map is fixed by the resolution of the raster grid. If we require a higher resolution, then we must recalculate the flow map on a raster grid with a higher resolution.
Our goal is to produce a traffic flow map from map matched routes, which can be utilised at any scale without recalculation, from the city level to the individual road level. In our proposed local alignment approach, we infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it. To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods. These include “node snapping” where the nearby boundary points of road segments are combined via statistical clustering, and “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments. We combine these novel algorithms into an iterative workflow which outputs a locally aligned flow map. This flow map is composed of road segments rather than raster grid cells, so its resolution is not limited by the raster grid resolution.
The outline of the paper is as follows. In Section 2, we construct an iterative workflow which computes a locally aligned flow map from noisy trajectories. In Section 3, we apply this workflow to synthetic trajectories from Los Angeles, USA and empirical trajectories from Hannover, Germany. We verify that, with a suitable tuning parameter choice, the locally aligned flow maps have high resolution and spatial accuracy. We then discuss some computational issues and end with some concluding remarks.
2 Local alignment of route segments for flow aggregation
We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point. We set $\mathcal{N}$ to be the OpenStreetMap (OSM) road network. Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry, following the Open Geospatial Consortium (OGC) terminology (OGC, 2010). We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$. This is a “multipoint” geometry (OGC, 2010). The points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$.
2.1 Map matching
We represent the output of a map matching algorithm $M$ from a trajectory $G$ as $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ which is a connected sequence of $n_{M}$ edges, i.e. a linestring geometry. In comparison to a trajectory $G$, all boundary points of the edges of a map matched route $M(G)$ are aligned to the road network. There is a large body of research on this difficult problem of map matching. We focus on the popular class of Hidden Markov Models (HMM) map matching algorithms. HMM methods iteratively build the map matched route by selecting the most likely next segment to connect to the current route using a probabilistic model. According to a review of map matching algorithms (Chao et al., 2020), HMM is a state-transition method. The other three classes are similarity, candidate-evolving, and scoring methods. Further details of these alternatives are found in Quddus et al. (2007), Chao et al. (2020). We leave this discussion here since our proposed methods are valid for any map matching algorithm, and concentrate on HMM map matching due to its accuracy and computational efficiency. Even if we restrict ourselves to HMM algorithms on the OSM road network, there are many off-the-shelf map matching APIs available. We focus on the Valhalla routing engine (https://valhalla.github.io/valhalla), which includes its highly recommended map matching API (Saki and Hagen, 2022).
To this off-the-shelf Valhalla map matching, we apply a post hoc modification to improve its suitability as input for flow map aggregation, as outlined in Algorithm 1. The inputs of ST_ROUTE are the trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding. These waypoints guide the route finding API to pass near them. Since we do not have an a priori single optimal value for the number of waypoints, we consider a range of $w$ values ${\bm{n}}_{W}=(n_{W,1},\dots,n_{W,w})$. In Step 1, we compute the initial map matched route $M(G)$ from the trajectory $G$ by calling the map matching API $M$. This initial map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points. In Steps 2–5, we loop over the number of waypoints in ${\bm{n}}_{W}$. In Steps 3–4, we take a sample of $n_{W,i}$ waypoints, where the first waypoint ${\bm{w}}_{1}$ is the start point of ${\bm{m}}_{1}$ and the $n_{W,i}$th waypoint is the end point of ${\bm{m}}_{n_{M}}$, and the intermediate waypoints ${\bm{w}}_{2},\dots,{\bm{w}}_{n_{W,i}-1}$ are sampled from the start points of the edges ${\bm{m}}_{2},\dots,{\bm{m}}_{n_{M}}$. In Step 5, we call the route finding API $R$ with the waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}}$. The result is the map matched route $R_{i}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})$. In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $R_{i}(G)$ and the trajectory $G$, for $i=1,\dots w$. The DTW distance is based on the lengths of all the distortions of $R_{i}(G)$ to achieve maximal alignment between $R_{i}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
Algorithm 1 ST_ROUTE – Map matched route
1:Input: $G$ trajectory, $M$ map matching API, $R$ route finding API, ${\bm{n}}_{W}$waypoints
2:Output: $R(G)$ map matched route
3:Compute initial map matched route $M(G)$ from empirical trajectory $G$
4:for $i:=1$ to $w$ do
5:     Set $n_{W}:=n_{W,i}$
6:     Sample waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from initial map matched $M(G)$
7:     Compute map matched routes $R_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
8:Select minimal route $R(G):=\operatorname{argmin}_{i\in\{1,\dots,w\}}\,\operatorname{DTW}(R_{i}(G),G)$
We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API. We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value. The map matched Hannover trajectories are displayed as the blue lines in Figure 1.
2.2 Flow aggregation
To the road network graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, we add the traffic flows on each of the network edges to form a flow map =$\mathcal{F}=\{(f,{\bm{f}}):{\bm{f}}\in\mathcal{N},f>0\}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ composed of $n_{\mathcal{F}}$ flows, where each ${\bm{f}}_{i}$ is composed of edges in $\mathcal{E}$, with traffic flow $f_{i}$, $i=1,\dots,n_{\mathcal{F}}$. Our goal is to estimate the traffic flows $f_{1},\dots,f_{n_{\mathcal{F}}}$ and the flow network graph $\mathcal{F}$. For exactly aligned trajectories, to compute the flow aggregation, we deploy the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018). That is, starting with the map matched routes =$\mathcal{R}=\{R(G_{1}),\dots,R(G_{n})\}$, the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$.
To compute an accurate flow map from noisy, empirical trajectories relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments. Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021). Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring. These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road level, and remains mostly suited to large-scale aggregations, such as desire line maps (Tobler, 1987). We do not consider edge bundling further since it is not suitable for our purposes. Rasterisation relies on converting the flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood. Whilst rasterisation is able to improve flow aggregation, it depends highly on the raster pixel neighbourhood size, and leads to a loss of spatial resolution.
We propose an alternative aggregation which solves both of these problems. We achieve this goal by local internal alignment between the map matched routes. By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph. By local alignment, we mean that we align segments of the routes, rather than the complete routes. Our proposal is divided into several algorithms, which we develop within the R statistical analysis environment due to its integrated access to advanced statistical and geospatial analysis methods.
2.3 Node snapping and node splitting
The proposed algorithm is ST_SNAPNODE, where the boundary points of the traffic flow linestrings are snapped to each other. Since the former are also nodes of the flow map, this gives the name to the algorithm. We focus on snapping these nodes since the linestring misalignments are in part caused by the existence of nodes which are close to each other but not exactly equal.
Since we are searching for points which are close to together, then this is well-suited to statistical clustering. There are many statistical clustering algorithms available, and we focus on hierarchical clustering. To resolve the computational bottleneck of hierarchical clustering, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering. We begin with an efficient single linkage clustering in the R package fastcluster (Müllner, 2013). Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains. In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so it is less likely to reach computational limits.
Algorithm 2 is a description of ST_SNAPNODE. The inputs are the $n_{\mathcal{F}}$ flow linestrings =$\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$. As its name suggests, all node points within $\varepsilon_{S}$ distance of each other are snapped together. In Step 1, we extract the boundary points of all flow linestrings. In Steps 2–3, we compute a single linkage clustering on all boundary points, and cut the dendrogram at height $\varepsilon_{S}$, resulting in $C^{\prime}$ clusters. In Steps 4–7, within each of these single linkage clusters, we compute a complete linkage clustering, and cut the dendrogram at height $\varepsilon_{S}$. This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other. In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the traffic flow values. In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters. In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid. In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$. In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
Algorithm 2 ST_SNAPNODE – Snap node clustering of linestrings
1:Input: $\mathcal{F}$ flow linestrings, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}^{*}$ snap node clustered flow linestrings
3:Extract boundary points $B(\mathcal{F}):=\{\operatorname{Start}({\bm{f}}_{1}),\operatorname{End}({\bm{f}}_{1}),\dots,\operatorname{Start}({\bm{f}}_{n_{\mathcal{F}}}),\operatorname{End}({\bm{f}}_{n_{\mathcal{F}}})\}$
4:Compute hierarchical clustering with single linkage on $B(\mathcal{F})$
5:Cut single linkage dendrogram at height $\varepsilon_{S}$ to compute $C^{\prime}$ cluster labels
6:for $i:=1$ to $C^{\prime}$ do
7:     Extract $i$th cluster of boundary points $B_{i}^{\prime}:=\{{\bm{b}}_{1}^{\prime},\dots,{\bm{b}}_{n_{B^{\prime}}}^{\prime}\}$
8:     Compute hierarchical clustering with complete linkage on $B_{i}^{\prime}$
9:     Cut complete linkage dendrogram at height $\varepsilon_{S}$ to compute $C^{\prime\prime}$ cluster labels
10:     for $j:=1$ to $C^{\prime\prime}$ do
11:         Extract $j$th cluster of boundary points $B_{j}^{\prime\prime}:=\{{\bm{b}}_{1}^{\prime\prime},\dots,{\bm{b}}_{n_{B^{\prime\prime}}}^{\prime\prime}\}$
12:         Extract corresponding flows $\{{\bm{f}}_{1}^{\prime\prime},\dots,{\bm{f}}_{n_{B^{\prime\prime}}}^{\prime\prime}\}$ from $\mathcal{F}$
13:         Compute ${\bm{b}}_{j}^{*}$ := weighted centroid of $\{{\bm{f}}_{1}^{\prime\prime},\dots,{\bm{f}}_{n_{B^{\prime\prime}}}^{\prime\prime}\}$, weights := $f_{1}^{\prime\prime},\dots,f_{n_{B^{\prime\prime}}}^{\prime\prime}$
14:Renumber collated complete linkage cluster labels to unique $C^{*}$ labels
15:for $i:=1$ to $C^{*}$ do
16:     Extract corresponding flow linestrings $\mathcal{F}_{i}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{B^{*}}}^{*},{\bm{f}}_{n_{B^{*}}}^{*})\}$ from $\mathcal{F}$
17:     Snap points of $\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ within $\varepsilon_{S}$ distance of cluster centroid to ${\bm{b}}^{*}_{i}$
18:     $\mathcal{F}_{i}^{*}$ := rejoin snapped and unsnapped points into linestrings
19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
20:Sort $\mathcal{F}^{*}$ by descending flow and length
Node snapping is focused near the boundary points of the flow linestrings, so it does not address misalignments far from the boundary points. Moreover there remain some linestrings whose intersection is not correctly computed by ST_OVERLINE_PLANR. The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
ST_SPLITNODE is the collation of a couple of algorithms which explicitly add the nodes at the intersections in the interior of linestrings, and splits these linestrings into simple linestring segments at these added nodes. This gives the name to this method. The inputs to Algorithm 3 are the flow linestrings $\mathcal{F}$ and the node splitting type $S$. Our method deploys for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2023), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2023). The first option tends to find fewer intersection nodes and is similar to the intersection computation performed by ST_OVERLINE_PLANR. The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation. We require both types of node splitting to compute a locally aligned flow map.
Algorithm 3 ST_SPLITNODE – Split nodes at interior intersections of linestrings
1:Input: $\mathcal{F}$ flow linestrings, $S$ node splitting type
2:Output: $\mathcal{F}^{*}$ node split flow linestrings
3:Initialise local network $\mathcal{N}^{*}$ with linestrings $\mathcal{F}$
4:if $S==$ “subdivision” then
5:     $\mathcal{F}^{*}:=$ sfnetworks::to_spatial_subdivision$(\mathcal{N}^{*})$
6:else if $S==$ “unary” then
7:     $\mathcal{F}^{*}:=$ geos::geos_unary_union$(\mathcal{N}^{*})$
In Figure 2(a) are the flow segments without any node splitting or node snapping. There are missing intersection nodes, and there are intersection nodes which are close to each other. In Figure 2(b) are the flow segments after node splitting ($S=$ “unary”) and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m). The result is that there are fewer flow segments with higher flow values. Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 2(b) is an improvement over Figure 2(a).
Refer to caption Refer to caption
Figure 2: Unsplit/unsnapped and node split/node snapped flow segments, with snap tolerance $\varepsilon_{S}=4$ m. (a) Unsplit and unsnapped. (b) Node split and node snapped.
2.4 Line blending to align similar linestrings with local reference
So far we have focused on improving the alignment of linestrings by resolving inconsistencies at their intersections. We now focus on aligning linestrings more generally. For this, we require a comparison relationship to establish an order of alignment of nearby linestrings. For two flow linestrings $(f_{1},{\bm{f}}_{1})$ and $(f_{2},{\bm{f}}_{2})$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon),\ f_{2}\leq f_{1}.$ (1)
The condition on the flow values $f_{2}\leq f_{1}$ means that we place a higher priority on the reference linestring ${\bm{f}}_{1}$ since its flow $f_{1}$ is larger than the flow $f_{2}$ on the candidate linestring ${\bm{f}}_{2}$. The buffer zone has flat edges, e.g. ST_BUFFER(endCapStyle="FLAT") in the R package sf (Pebesma, 2018), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$. This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence, to be candidates to be aligned to ${\bm{f}}_{1}$.
Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$. The set of $n_{\mathcal{F}_{\rm cand}}$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set. We call our approach “line blending”, since we blend the candidate linestrings onto the reference linestring. The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$. The output are the modified reference linestring and $n_{\mathcal{F}_{\rm cand}}$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments. In Step 1, we initialize a local network graph $\mathcal{N}^{*}$ with the reference linestring. In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023). ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$. The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring. In Steps 4–5, we extract the $n_{\mathcal{F}_{\rm cand}}$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$. Step 6 involves collating the blended candidate linestrings. Step 7 is the equivalent for the blended reference linestring.
Algorithm 4 ST_LINEBLEND – Blend candidate linestrings onto reference linestring
1:Input: ${\bm{f}}_{\rm ref}$ reference, $\mathcal{F}_{\rm cand}$ candidate flow linestrings, $\varepsilon$ blend tolerance
2:Output: ${\bm{f}}^{*}_{\rm ref}$ blended reference, $\mathcal{F}^{*}_{\rm cand}$ blended candidate flow linestrings
3:Initialise local network graph $\mathcal{N}^{*}$ with ${\bm{f}}_{\rm ref}$
4:Extract all points $G_{\rm cand}$ of candidate linestrings and flow values $f_{\rm cand}$ from $\mathcal{F}_{\rm cand}$
5:Update network by blending candidate points $\mathcal{N}^{*}:={\tt ST\_NETWORK\_BLEND}(\mathcal{N}^{*},G_{\rm cand},\varepsilon)$
6:for $i:=1$ to $n_{\mathcal{F}_{\rm cand}}$ do
7:     ${\bm{f}}^{*}_{{\rm cand},i}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm cand},i})$ to $\operatorname{End}({\bm{f}}_{{\rm cand},i})$ along network $\mathcal{N}^{*}$
8:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
9:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points. If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values. So we assign the rounded value of the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
In Figure 3(a), we illustrate this line blending with the flow linestrings $(7,ABC)$ in blue and $(2,AD)$ in orange. As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them. Since $AD\subset{\tt ST\_BUFFER}(ABC,4\,\mathrm{m})$ (the pale blue rectangle), and $2<7$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1). Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring. We also add $D^{\prime}$ to the reference linestring. The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,ABD^{\prime}C)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,ABD^{\prime})$. Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $ABD^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,ABD^{\prime})$ and $(2,D^{\prime}C)$. We take the rounded value of the weighted mean of these two flow linestrings. The result is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 3(b).
Refer to caption Refer to caption
Figure 3: Line blending leads to correct flow aggregation. (a) Before line blending and flow aggregation. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange. (b) After blending candidate into reference linestring, with tolerance $\varepsilon=4$ m, and flow aggregation. Blended linestring is $(9,ABD^{\prime}C)$.
After aggregation, the projected point $D^{\prime}$ in $ABD^{\prime}C$ appears to be redundant. Suppose that we have another linestring $(3,ABC)$ without $D^{\prime}$. Then ST_OVERLINE_PLANR does not aggregate $(9,ABD^{\prime}C)$ with $(3,ABC)$ even if $D^{\prime}$ is exactly coincident on the segment $BC$, since this algorithm does not explicitly add any missing intersection nodes prior to aggregation (Morgan and Lovelace, 2021). It outputs three linestrings $(12,AB),(9,BC),(3,BD^{\prime}C)$, rather than a single linestring $(12,ABD^{\prime}C)$. So we retain these intermediate points in the flow linestrings to assist in subsequent flow aggregation.
More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending. Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map. This procedure ST_SNAP_CAND_TOUCH is outlined in Algorithm 5. The inputs are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$. In Steps 1–8, we iterate over each candidate-touching linestring. In Steps 2–3, we compute the intersections between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances. In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point. In Step 8, if the intersection point is not within $\varepsilon_{S}$ distance, then we snap it to the nearest point on ${\bm{f}}_{\rm ref}$. These snapping operations are similar to those in Steps 15–16 in ST_SNAPNODE in Algorithm 2, and ensure that we maintain connectivity between $\mathcal{F}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$. In Step 9, we collate these snapped linestrings.
Algorithm 5 ST_SNAP_CAND_TOUCH – Snap candidate-touching linestrings onto reference
1:Input: ${\bm{f}}_{\rm ref}$ reference, $\mathcal{F}_{\rm cand}$ candidate, $\mathcal{F}_{\rm ct}$ candidate-touching flow linestrings, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}^{*}_{\rm ct}$ snapped candidate-touching flow linestrings
3:for $i:=1$ to $n_{\mathcal{F}_{\rm ct}}$ do
4:     ${\bm{g}}_{{\rm ct},i}^{*}:={\bm{f}}_{{\rm ct},i}\cap\{\operatorname{Start}(\mathcal{F}_{{\rm cand}}),\operatorname{End}(\mathcal{F}_{{\rm cand}})\}$
5:     $d_{i,\operatorname{Start}}^{*}:={\tt ST\_DIST}({\bm{g}}_{{\rm ct},i}^{*},\operatorname{Start}({\bm{f}}_{\rm ref}))$; $d_{i,\operatorname{End}}^{*}:={\tt ST\_DIST}({\bm{g}}_{{\rm ct},i}^{*},\operatorname{End}({\bm{f}}_{\rm ref}))$
6:     if ($d_{i,\operatorname{Start}}^{*}\leq\varepsilon_{S}$ and $d_{i,\operatorname{Start}}^{*}\leq d_{i,\operatorname{End}}^{*}$then
7:         ${\bm{f}}^{*}_{{\rm ct},i}:=$ snap ${\bm{f}}_{{\rm ct},i}$ to $\operatorname{Start}({\bm{f}}_{\rm ref})$
8:     else if ($d_{i,\operatorname{End}}^{*}\leq\varepsilon_{S}$ and $d_{i,\operatorname{End}}^{*}\leq d_{i,\operatorname{Start}}^{*}$then
9:         ${\bm{f}}^{*}_{{\rm ct},i}:=$ snap ${\bm{f}}_{{\rm ct},i}$ to $\operatorname{End}({\bm{f}}_{\rm ref})$
10:     else ${\bm{f}}^{*}_{{\rm ct},i}:=$ snap ${\bm{f}}_{{\rm ct},i}$ to nearest point of ${\bm{f}}_{\rm ref}$
11:Collate $\mathcal{F}^{*}_{\rm ct}:=\{{\bm{f}}^{*}_{{\rm ct},1},\dots,{\bm{f}}^{*}_{{\rm ct},n_{\mathcal{F}_{\rm ct}}}\}$
Figure 4(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $(7,ABC)$ (blue), the candidate linestring $(2,AD)$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue). The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$. The orange line is entirely within the pale blue buffer zone, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring. If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e., ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$. If we apply ST_SNAP_CAND_TOUCH with snap tolerance $\varepsilon_{S}=4$ m to the candidate-touching linestring, then the intersection point $D$ is projected to the boundary point $C$, i.e., ${\bm{f}}^{*}_{\rm ct}=(5,CE)$. Thus line blending, candidate-touching snapping, and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
Refer to caption Refer to caption
Figure 4: Snap candidate-touching linestrings to reference linestring after line blending, with blend tolerance $\varepsilon=4$ m, snap tolerance $\varepsilon_{S}=4$ m. (a) Before line blending and snapping. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ orange, and candidate touching linestring $(5,DE)$ purple. (b) After line blending and snapping. Reference linestring becomes $(9,ABD^{\prime}C)$, and candidate-touching linestring $(5,CE)$.
We generalise the reference linestring to be a sequence of $k$ connected edges, each with their own flow. If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings. Due to computational limitations, we retain that candidates be single edges with single flow values. The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
We have described blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges). The next step is to determine the priority for line blending for a set of reference linestrings. We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only. These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them. The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except when $f_{1}=f_{2}$. In this case, then we designate the longer linestring to be the reference.
The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$. The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings. In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e., all paths composed of connected, unique edges) with $k$-edges. In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path. The function ${\rm Edge}(\cdot)$ extracts the component edges from a $k$-edge path. In Step 5, we sort the flow linestrings, in descending order of their flow values and length. In Steps 6–8, we construct a maximal set of non-overlapping $k$-edge paths $\mathcal{K}^{*}$. We initialise $\mathcal{K}^{*}$ to contain the first path. We iterate through $\mathcal{K}$ and, if this path in $\mathcal{K}$ does not overlap the current $\mathcal{K}^{*}$, then we add it to $\mathcal{K}^{*}$. For $k=1$, we bypass Steps 2–8. In Steps 9–18, we step through the ${\bm{k}}_{i}^{*}\in\mathcal{K}^{*}$, starting from the linestring with the highest flow and length. In Steps 11–13, if the current flow linestring ${\bm{k}}_{i}^{*}$ has no incident edges in the candidate linestrings $\mathcal{F}_{\rm cand}$, then we set it to be a reference linestring. In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1). In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$, and $\mathcal{F}_{\rm ct}$. In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets.
Algorithm 6 ST_LINEBLEND_PRIORITY – Compute line blending priority
1:Input: $\mathcal{F}$ flow linestrings, $k$edges, $\varepsilon$ blend tolerance
2:Output: $(\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct})$ reference, candidate, candidate-touching linestrings
3:Initialise network graph $\mathcal{N}^{*}$ from linestrings $\mathcal{F}$
4:Extract all simple paths of length $k$ edges $\mathcal{K}:=\{{\bm{k}}_{1},\dots,{\bm{k}}_{n_{\mathcal{K}}}\}$ from $\mathcal{N}^{*}$
5:for $i:=1$ to $n_{\mathcal{K}}$ do
6:     $f_{i}$ := weighted mean of $k$ flows, weight := len(Edge(${\bm{k}}_{i}$))
7:Sort $\mathcal{K}$ by descending flow and combined length
8:Initialise $\mathcal{K}^{*}:=\{{\bm{k}}_{1}\}$
9:for $i:=2$ to $n_{\mathcal{K}}$ do
10:     if (${\bm{k}}_{i}\cap\mathcal{K}^{*}=\{\}$then $\mathcal{K}^{*}:=\mathcal{K}^{*}\cup\{{\bm{k}}_{i}\}$
11:Initialise $\mathcal{Q}:=\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm ct}:=\{\}$
12:for $i:=1$ to $n_{\mathcal{K}^{*}}$ do
13:     if (${\rm Edge}({\bm{k}}_{i}^{*})\notin\mathcal{F}_{\rm cand}$then
14:         Set reference linestring ${\bm{k}}_{{\rm ref}}:={\bm{k}}_{i}^{*}$
15:         Search $\mathcal{F}_{{\rm cand}}^{*}:=\{{\bm{f}}\in\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\{{\bm{k}}_{\rm ref}\}\cup\mathcal{F}_{\rm cand})\colon{\bm{f}}\subseteq{\tt ST\_BUFFER}({\bm{k}}_{\rm ref},\varepsilon)\}$
16:         if $\mathcal{F}_{\rm cand}^{*}\neq\{\}$ then
17:              Update $\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm ref}\cup\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm cand}^{*}$
18:              Search $\mathcal{F}_{{\rm ct}}:=\{{\bm{f}}\in\mathcal{F}\backslash\{{\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\}\colon{\tt ST\_TOUCHES}(\mathcal{F}_{{\rm cand}}^{*},{\bm{f}})\}$
19:              Update $\mathcal{Q}:=\mathcal{Q}\cup\{(\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}^{*},\mathcal{F}_{\rm ct}^{*})\}$
20:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
The local alignment of flow segments, in Algorithm 7, is established by combining Algorithms 26. The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges $k$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$. In Step 1, we node split the flow linestrings $\mathcal{F}$. In Step 2, we node snap the flow linestrings. In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$. In Steps 5–8, we iterate over each reference linestring. In Step 6, we update the reference and candidate flow linestrings by blending the candidate linestrings. In Step 7, we apply ST_OVERLINE_PLANR within each of the $k$-edges of the updated reference linestrings. In Step 8, we compute the weighted mean flow, weighted by the length of the corresponding sub-segments, and assign it as the flow value to all sub-segments. The result from Steps 6–8 is a modified $k$-edge reference linestring with $k$ aggregated flows. In Step 9, we snap the candidate-touching linestrings to the reference linestring to maintain connectivity. In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.
Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align segment flows using line blending
1:Input: $\mathcal{F}$ flow linestrings, $S$ split node, $k$edges, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}^{*}$ aggregated aligned flow linestrings
3:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
4:$\mathcal{F}:={\tt ST\_SNAPNODE}(\mathcal{F},\varepsilon_{S})$
5:$\{\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}\}:={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
6:$\Edgesubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
7:for $i:=1$ to $n_{\mathcal{F}_{\rm ref}}$ do
8:     $\{{\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i}^{*}\}:={\tt ST\_LINEBLEND}({\bm{f}}_{{\rm ref},i},\mathcal{F}_{{\rm cand},i},\varepsilon)$
9:     ${\bm{f}}_{{\rm ref},i}^{*}:={\tt ST\_OVERLINE\_PLANR}({\bm{f}}_{{\rm ref},i}^{*}\cup\mathcal{F}_{{\rm cand},i}^{*})$
10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edge}({\bm{f}}_{{\rm ref},i}^{*}))$
11:     $\mathcal{F}_{{\rm ct},i}^{*}:={\tt ST\_SNAP\_CAND\_TOUCH}({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i},\mathcal{F}_{{\rm ct},i},\varepsilon_{S})$
12:$\mathcal{F}^{*}:=\mathcal{F}_{\rm ref}^{*}\cup\mathcal{F}_{\rm ct}^{*}\cup\mathcal{F}^{c}$
13:Sort $\mathcal{F}^{*}$ in descending order of flow and length
In Figure 5(a) are the flow segments without any line blending, where, for example, the 3 segments with flow 2, 152 and 163, should be blended into a single segment. In Figure 5(b) are the flow segments after line blending. These 3 separate segments are blended into a segment with flow 317 (pink), which is the sum of their individual flows. In contrast, the segments with flows 8, 12 and 16, forming part of the T-junction, are blended into a segment with a weighted mean of 15. Due to the action of ST_OVERLINE_LINEBLEND, the line blended flow segments in Figure 5(b) are an improvement over Figure 5(a).
Refer to caption Refer to caption
Figure 5: (a) Flow segments without line blending. (b) Flow segments with line blending. Label is traffic flow.
To compute a locally aligned flow map, we combine Algorithms 17 into an initial and an iterative phase. The initial phase ST_OVERLINE_INITIAL is described in Algorithm 8. Its inputs are the map matched routes $\mathcal{R}$, the split node type $S$, and the simplify tolerance $\varepsilon_{D}$. In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce a flow map $\mathcal{F}$. In Step 2, we apply ST_SPLITNODE to this flow map. In Step 3, we employ the standard ST_SIMPLIFY with simplify tolerance $\varepsilon_{D}$. These simplified linestrings are modified so that all modified segments are at most $\varepsilon_{D}$ distance from the unmodified segments (Ramer, 1972, Douglas and Peucker, 1973). These simplified linestrings usually lead to more overlapping segments, which assist the flow aggregation in Step 4.
Algorithm 8 ST_OVERLINE_INITIAL – Compute initial locally aligned flow map
1:Input: $\mathcal{R}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance
2:Output: $\mathcal{F}$ initial aligned flow map
3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$
4:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
5:$\mathcal{F}:={\tt ST\_SIMPLIFY}(\mathcal{F},\varepsilon_{D})$
6:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
With an initial flow map on hand, we proceed iteratively towards a locally aligned flow map in Algorithm 9. The inputs of ST_OVERLINE are the initial flow map $\mathcal{F}$, the split node type $S$, the blend tolerance $\varepsilon$, the snap tolerance $\varepsilon_{S}$, the number of connected edges ${\bm{k}}$, and the maximum number of iterations $j_{\max}$. Steps 1–6 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings. For each $k$ in ${\bm{k}}$, we iterate until $j_{\max}$ is reached or two consecutive flow maps are identical. Within each iteration, we search for the $k$-edge reference linestrings, blend the candidate linestrings, and snap the candidate-touching linestrings, and compute the flow aggregation. In Steps 7–10 is some housekeeping in ST_OVERLINE_PRUNE, where we remove pseudo nodes, remove leaf edges with low flow values, and replace thin loops with non-looped linestrings.
Algorithm 9 ST_OVERLINE – Compute iterated locally aligned flow map
1:Input: $\mathcal{F}$ initial flow map, $S$ split node, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
2:for $k$ in ${\bm{k}}$ do
3:     $\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
4:     while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
5:         $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
6:         $\mathcal{F}:={\tt ST\_OVERLINE\_LINEBLEND}(\mathcal{F},k,\varepsilon,\varepsilon_{S})$
7:         $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
8:$\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
9:while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
10:     $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
11:     $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
2.5 Workflow
Our proposed workflow from an input of trajectory data to a locally aligned flow map is illustrated in Figure 6. We begin with the trajectories $\mathcal{G}$. If the measurement errors in these trajectories are small enough to be input directly into ST_OVERLINE, then we set $\mathcal{R}=\mathcal{G}$. Otherwise to obtain $\mathcal{R}$, we carry out map matching using ST_ROUTE for each number of waypoints ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value. For the flow map aggregation, we begin with an iteration of ST_OVERLINE_INIITAL with simplify tolerance $\varepsilon_{D}=1$ m to create an initial flow map. In Figure 6, the dotted rectangle surrounding ST_LINEBLEND_PLANR and ST_SNAPNODE indicates that these algorithms are inputs to ST_OVERLINE_INIITAL. For the subsequent steps, we set the maximum number of iterations $j_{\max}=20$, and the snap tolerance $\varepsilon_{S}=\varepsilon$. We call ST_OVERLINE with #edges ${\bm{k}}=1,2$, and node split type $S=$ “subdivision”. At this early stage, the search for connected $k$-edges with $k>2$ can be computationally intensive, and “subdivision” node splitting provides more stable line blending priority. The dotted rectangle indicates the algorithms which are inputs to ST_OVERLINE. We follow with a call to ST_OVERLINE with “unary” node splitting. The “unary” node splitting is usually applied after “subdivision” node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority. After these two steps, the flow map has begun to stabilise, and 3-, 4-edge searches become computationally feasible with the lower number of flow segments. We iterate ST_OVERLINE with ${\bm{k}}=1,2,3,4,S=$ “unary”, until the flow map $\mathcal{F}$ converges. In this workflow, the only unspecified tuning parameter is the blend tolerance $\varepsilon$, whose properties we investigate in the next section.
Figure 6: Workflow for locally aligned flow map. Input are trajectories $\mathcal{G}$, output is locally aligned flow map $\mathcal{F}$. Blend tolerance $\varepsilon$ is tuning parameter.
3 Results and discussion
3.1 Synthetic trajectories
Despite the ubiquity of GNSS-enabled devices, there is a lack of gold standard empirical trajectories and flow maps, due to privacy concerns and other reasons. Our solution is to rely on noise-free synthetic trajectories which are exactly aligned to a theoretical road network (Anastasiou et al., 2022a). These authors generate synthetic trajectories from an origin-destination matrix, and make available a set of 1.5 million synthetic trajectories for the metropolitan area of Los Angeles, USA (Anastasiou et al., 2022b). For illustrative purposes, we focus on a sample of $n=100$ trajectories, in EPSG 6339 (NAD83 (2011)/UTM zone 10N) coordinates, in Culver City (located about 16 km from downtown Los Angeles) which pass near the intersection of Main Street and Venice Boulevard, as indicated by the solid yellow line in Figure 7. We add a random uniform noise of ±$\pm 5$ m to the noise-free trajectories to obtain the noisy trajectories, displayed as the green lines in Figures 7(a–b). This noise mimics the typical measurement error of 5 m observed in empirical trajectories from GNSS-enabled smartphones (Van Diggelen and Enge, 2015). The reference noise-free flow map is in Figure 7(c), and the rasterised flow map in Figure 7(d). Morgan and Lovelace (2021) do not recommend a value for tuning the rasterisation grid resolution, so we select $\varepsilon_{R}=2$ m subjectively so that the raster grid cells do not exceed the road widths.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Synthetic Culver City trajectories $(n=100)$. (a) Noisy trajectories (green lines). (b) Zoom of noisy trajectories near intersection of Main St and Venice Bd (solid yellow line). (c) Reference noise-free flow map. (d) Rasterised flow map with $\varepsilon_{R}=2$ m. Label in flow maps is traffic flow.
For our locally aligned flow maps, the free tuning parameter is the blend tolerance parameter $\varepsilon$. So we investigate the effect of $\varepsilon$ on the accuracy of flow aggregation. For the blend tolerance $\varepsilon=1$ m, the flow map in Figure 8(a) is unable to consistently blend the map matched routes correctly since 1 m is typically smaller than the remaining discrepancies after map matching in ST_ROUTE. For the blend tolerance $\varepsilon=5$ m, the central road segments labelled with flow 51 and 45 in Figure 8(b), which are close to 49 and 45 in the reference flow map. For the blend tolerance $\varepsilon=10$ m, the central road segments are also labelled with flow 50 and 45 in Figure 8(c), similar to with blend tolerance $\varepsilon=5$ m in Figure 8(b). Near the intersection of Venice Bd and Main St (solid yellow line), the flow segment (with flow 17) is poorly aligned to the road network, unlike the more accurate alignment in Figure 8(b).
Refer to caption Refer to caption
Figure 8: Effect of blend tolerance $\varepsilon$ on synthetic Culver City flow maps. (a) Flow map with $\varepsilon=1$ m. (b) Flow map with $\varepsilon=5$ m. (c) Flow map with $\varepsilon=10$ m. Label is traffic flow.
To complement the above heuristic comparisons, we define a quantitative discrepancy between a flow map $\mathcal{F}$ and the reference flow map $\mathcal{F}_{\rm ref}$. First, we compute the transect points for each linestring ${\bm{f}}$ from $\mathcal{F}$: if $\operatorname{len}({\bm{f}})>\varepsilon_{T}$ m, then a point is placed at every $\varepsilon_{T}$ m of ${\bm{f}}$; or if $\operatorname{len}({\bm{f}})\leq\varepsilon_{T}$ m, then a single point is placed at the centroid of ${\bm{f}}$. So $\mathcal{T}=\{{\bm{t}}_{1},\dots,{\bm{t}}_{n_{\mathcal{T}}}\}$ is a set of $n_{\mathcal{T}}$ transect points. For each ${\bm{t}}_{i},i=1,\dots,n_{\mathcal{T}}$, we find the closest linestring in $\mathcal{F}_{\rm ref}$, that is, ${\bm{f}}_{{\rm ref},i}^{*}=\operatorname{argmin}_{{\bm{f}}_{\rm ref}\in\mathcal{F}_{\rm ref}}\mathrm{dist}({\bm{t}}_{i},{\bm{f}}_{\rm ref})$. The flow errors are $\operatorname{Err}_{i}=|f_{i}-f_{{\rm ref},i}^{*}|$ for the transect points on ${\bm{t}}_{i}\in\mathcal{T}\subseteq\mathcal{F}$. We reverse the order of the comparison to obtain the flow errors $\operatorname{Err}_{j}=|f_{{\rm ref},j}-f_{j}^{*}|$ for the transect points on ${\bm{t}}_{j}\in\mathcal{T}_{\rm ref}\subseteq\mathcal{F}_{\rm ref}$, and concatenate these two sets of flow errors.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Flow errors for synthetic Culver City flow maps. (a) Flow map with $\varepsilon=1$ m. (b) Flow map with $\varepsilon=5$ m. (c) Flow map with $\varepsilon=10$ m. (d) Rasterised flow map with $\varepsilon_{R}=2$ m. Flow segments are coloured lines, transect points are grey circles, flow errors are labels.
In Figure 9(a), we display the transect points with $\varepsilon_{T}=50$ m, as the grey circles, superposed over the flow map $\mathcal{F}$. Figures 9(a–c) are the flow errors for the locally aligned flow maps with blend tolerance $\varepsilon=1,5,10$ m, and Figure 9(d) is for the rasterised flow map. The $\varepsilon=1$ m flow map and the rasterised flow map lead to many high error values. Whereas for the other two locally aligned flow maps, the flow errors are similarly low.
We repeat the flow map error calculations for $N=10$ re-samples of $n=100$ Culver City trajectories. The mean and SD of the flow errors for the locally aligned flow maps are $3.061\pm 0.565$ ($\varepsilon=1$ m), $0.446\pm 0.047$ ($\varepsilon=5$ m), $0.477\pm 0.048$ ($\varepsilon=10$ m), and for the rasterised flow map $1.514\pm 0.201$ ($\varepsilon_{R}=2$ m). Setting $\varepsilon=1$ m leads to under-blending since $\varepsilon$ is smaller than the misalignments in the map matched routes, and setting $\varepsilon=10$ m leads to over-blending since $\varepsilon$ is larger than the distance between separate road segments. The $\varepsilon_{R}=2$ m rasterised flow map is similarly under-blended like the $\varepsilon=1$ m flow map. So we tentatively recommend a value of about $\varepsilon=5$ m for the blend tolerance parameter since it leads to optimal line blending.
3.2 Empirical trajectories
We return to the empirical Hannover trajectories. From the 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m, and follow the workflow in Figure 6. We map match these 1177 trajectories with ST_ROUTE, and we continue with 1152 of them which have a sufficiently high quality match, where the Hausdorff distance $\operatorname{dist}_{\operatorname{Haus}}(R(G),G)<100$ m, and the ratio $\operatorname{len}(R(G))/\operatorname{len}(G)<1.1$, for a trajectory $G$ and a map matched route $R(G)$. That is, we discard 25 trajectories or a 2.02% failure rate for the Valhalla map matching, which is consistent with previous studies (Saki and Hagen, 2022).
After some trial and error, we set the blend tolerance of $\varepsilon=4$ m, which is 1 m less than the 5 m tentatively recommended from the synthetic trajectories computations above. The workflow leads to a locally aligned flow map displayed in Figure 10(a). The corresponding rasterised flow map is in Figure 10(b). At the city level, both of these flow maps visualise well the traffic flows. Though if we zoom into Klusreide (solid yellow circle), then we observe that in Figure 10(c) the locally aligned map displays the aggregated flows on linear road segments in comparison to the pixelated road segments of the rasterised flow map in Figure 10(d).
Whilst we do not have a reference flow map for these empirical trajectories, like we had for the synthetic Los Angeles trajectories, we are still able to compute a proxy flows based on empirical counts of the map matched routes with orthogonal line transects from the flow map road segments. Along Klusriede, the locally aligned flow map is composed to two road segments (in pink) with 0 and 2 flow errors) corresponding to the proxy flows 312 and 315. For the rasterised flow map, some grid cells in the pink region have flow error 2. Though rasterisation some times leads to double counting and thus to much larger flow errors, e.g. in the orange region near the T-junction, the estimated flow is 643 whereas the proxy flow is 315, so the flow error is 328. The locally aligned map does not suffer from these rasterisation effects, so its flow aggregation is more accurate.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 10: Hannover flow maps. (a) Locally aligned flow map at city level. (b) Rasterised flow map at city level. Solid yellow circle is Klusriede. (c) Locally aligned flow map at road segment level. (d) Rasterised flow map at road segment level. Colour scale is traffic flow, label is traffic flow error.
Our future goal is the integration of the local alignment algorithms into an industry standard web GIS. In the mean time, we make available interactive web maps at https://mvstat.net/traj which allow the user to freely explore our proposed multi-scale traffic flow maps from the individual road to the city-wide level.
3.3 Computation
We report the execution times for the flow map computations for the synthetic and empirical trajectories. Since it is difficult to reliably reproduce execution times on different systems, we provide our execution times as a guide only. These are obtained on a stand-alone set-up running Ubuntu 24.04 with 16Gb RAM and 4 × Intel i5 CPU @ 3.10 GHz. We deploy a local dockerised Valhalla routing engine API (https://github.com/gis-ops/docker-valhalla) (GIS OPS, 2023) rather than relying on a web-based service. Since the local map matching API requests can be efficiently parallelised on a stand-alone machine, we are confident that ST_ROUTE is suitably optimised.
For $n=100$ synthetic Culver City trajectories, ST_ROUTE executes in 188 s. For $n=1000$ synthetic trajectories, it executes in 2233 s, which is 10 times longer for a 10-fold increase in $n$. So it is straightforward to infer a precise execution time of ST_ROUTE for large sample sizes from a small sample size. The execution time for ST_OVERLINE increases as $n$ increases, but not in the same uniform manner as ST_ROUTE as the former also depends on the value of $\varepsilon$. For $n=100$, the execution times for ST_OVERLINE are 1777 s ($\varepsilon=1$ m), 1049 s ($\varepsilon=5$ m), 880 s ($\varepsilon=10$ m); and for $n=1000$, they are 15 457 s ($\varepsilon=1$ m), 6002 s ($\varepsilon=5$ m), 3897 s ($\varepsilon=10$ m). So it is difficult to infer precisely the execution times for ST_OVERLINE for larger sample sizes and smaller blend tolerances from smaller $n$ and larger $\varepsilon$. This is because the line blending (ST_LINEBLEND) is parallelised, but the line blending priority (ST_LINEBLEND_PRIORITY) remains a serial computation.
With an optimal value of the blend tolerance $\varepsilon=5$ m, the total execution time of the workflow for the $n=100$ Culver City trajectories is 1238 s or $\sim$21 mins, and for $n=1000$ is 8236 s or $\sim$137 mins. The total execution time for the $n=1183$ Hannover trajectories with $\varepsilon=4$ m is 5545 s or $\sim$92 mins. Thus locally aligned flow maps are currently suitable for offline analyses of small and medium sized data sets. Due to time and computing constraints, (a) data sets with a larger number of trajectories (like the full set of 1.5 million synthetic Los Angeles trajectories) and (b) online computations were out of scope. For these to be computationally feasible, we require additional implementation improvements in the local alignment algorithms, which we leave for future development.
4 Conclusion
We have introduced an iterative workflow which takes noisy GNSS vehicle trajectories as input and gives an accurate traffic flow map as output. Local alignment of flow segments is the key innovation. To carry out this local alignment, we defined a spatial relation to set up local reference flow segments, which then allowed us to align other nearby flow segments to this local reference segment. We presented solid evidence for the high level of spatial resolution and accuracy for our proposed locally aligned flow map for both synthetic and empirical trajectories. We demonstrated that they effectively reduce the noise of the input trajectories, and accurately aggregate the traffic flows on all road segments, and so they are a valuable addition to the toolkit for multi-scale traffic flow analysis.